Find the scalar x- and y-components of the following displacements in the XY-plane: (a) 300 cm at 127° and (b) 500 cm at 220°


1
Expert's answer
2022-03-25T03:23:42-0400

(a) "x=300\sin{127^0}=239.59 cm."

"y=300\cos{127^0}=-180.54 cm."


(b) "x=500\sin{220^0}=-321.39cm."

"y=500\cos{220^0}=-383.02 cm."